What is the molarity of a HCl solution if it takes 15.6 mL of a 1.00 M NaOH solution to titrate 25.0 mL of the acid to a neutral endpoint?


1
Expert's answer
2021-05-11T05:57:06-0400

Molarity of "HCl" can be calculated as,


"\dfrac{x \hspace{2mm}mol \hspace{2mm}of\hspace{2mm} HCl}{1L \hspace{2mm}of\hspace{2mm} HCl\hspace{2mm} solution} = \dfrac{25}{15.6}(\dfrac{10^3}{1L})(\dfrac{0.115\hspace{2mm}mol\hspace{2mm} NaOH}{10^3\hspace{2mm}mL\hspace{2mm} NaOH\hspace{2mm}solution})(\dfrac{1 mol\hspace{2mm} HNO_3}{1 mol\hspace{2mm} NaOH})"


From this equation we can calculate the molarity of "HCl" .